if an artery took up one fourth of a 25 mm field of view, how big was the artery
Answer #1

The field of view is = 25 mm

The area occupied by artery = 1/4 of total area

So the size of artery = (1/4)( Field of view)

= 1/4 x 25 mm

= 6.25 mm

Hence the size of artery is 6.25 mm.

It was a pleasure answering your question. Hope this helps...

Good luck!